I need to know the method of finding the answer the following problem.
Jim can fill a pool carrying buckets of water in 30 minutes. Sue can do the same job in 45 minutes. Tony can do the same job in 1 1/2 hours. How quickly can all three fill the pool together?

Answer by checkley77(12844)

x/30+x/45+x/90=1
(3x+2x+1x)/90=1
6x/90=1
6x=90
x=15 min. to fill the pool when they work together.